OpenStudy (anonymous):

Name the property illustrated by the statement: If -11a + 2 = -3a, then -3a = -11a + 2.

jimthompson5910 (jim_thompson5910):

Check out this page: http://library.thinkquest.org/2647/algebra/post.htm

jimthompson5910 (jim_thompson5910):

It lists the rules of algebra.

OpenStudy (anonymous):

Oh alright, So would be it Additive Property?

jimthompson5910 (jim_thompson5910):

no, the additive property is "if a = b, then a + c = b + c" (listed in second section of the page)

jimthompson5910 (jim_thompson5910):

have a second look at the page

OpenStudy (anonymous):

Oh so could you tell me what it would be?

jimthompson5910 (jim_thompson5910):

which equation/property best matches what you have

OpenStudy (anonymous):

Or wait would it be the Subtractive Property?

jimthompson5910 (jim_thompson5910):

Subtraction Property of Equality: if a = b, then a - c = b - c so no

jimthompson5910 (jim_thompson5910):

do you see why these aren't the answers?

OpenStudy (anonymous):

Hummm, Yes somewhat. So our answers could be Reflexive, Symmetric, Transitive, or Subsitution?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

compare your original equation with the equations they have on that page

OpenStudy (anonymous):

Hummm alright so would it be a Symmetric Property?

jimthompson5910 (jim_thompson5910):

yes it would be, the symmetric property basically says that if two numbers are equal, then you can write them as first number = second number or second number = first number the order doesn't matter